OpenStudy (anonymous):

1 and 2 are supplementary angles . m 1=x-39, and m 2 =x+61. find the measure of each angle.

OpenStudy (anonymous):

supplementary means they add up to 180 degrees. so m1 + m2 = 180. substitute what 1 and 2 equal for m1 and m2. (x-39) + (x + 61) = 180. x + x -39 + 61 = 180. 2x + 22 = 180 2x = 158 x = 79 Now find m1 and m2. m1 = x -39, or 79-39 m1 = 40. m2 = x + 61, or 79 + 61 m2 = 140. 40 + 140 equals 180, so they're supplementary.